OpenStudy (anonymous):

Please help me understand how to do this problem and simplify. 1/x+2 + 1/x-2

OpenStudy (anonymous):

Have a common denominator

OpenStudy (anonymous):

\[ \frac{1}{x+2}+\frac{1}{x-2}=\frac{x-2}{x^2-4}+\frac{x+2}{x^2-4} \] Can you finish it now?

OpenStudy (anonymous):

\[ \frac{1}{x+2}+\frac{1}{x-2}=\frac{x-2}{x^2-4}+\frac{x+2}{x^2-4}=\frac{2 x}{x^2 -4} \]